### v1

@@ 5.2 Covariance estimation and scale cuts
The work in [100] investigated in detail how to efficiently extract the cosmological signal from MeerKLASS data and also estimate the errors on any measurements. The far-left panel of Figure 4 again shows the cross-correlation power spectrum between the MeerKLASS deep field and the GAMA galaxies but, instead of the 1D spherically-averaged power, this is cylindrically averaged into ($_{\perp}, kk_{\parallel}$) bins. This gives a more detailed window into where the signal exists in this analysis. The panel clearly shows the strong effect of the instrumental beam, where the power on small transverse scales (large $k_{\perp}$) is suppressed to nearly zero. At the opposite end, the strength of the signal at extremely small $k$ is also unreliable, which can be attributed both to survey geometry limitations and to residual contamination from foregrounds. This interpretation is supported by the transfer function analysis (shown in the second panel), which indicates that nearly all of the cosmological signal is lost ($\mathcal{T}\,{\sim}\,0$) at the very smallest $k$-modes owing to their strong degeneracy with foreground spatial structure.
@@
A major part of the investigation in [100] was to explore the use of the transfer function not only as a correction for signal loss but also as a means of estimating the covariance of the power spectrum. In earlier analyses [e.g. 56], analytical error estimates were adopted, assuming that thermal noise dominated the error budget and that Gaussian statistics applied. However, with the improved sensitivity of MeerKLASS, these assumptions no longer hold: thermal noise is subdominant on many scales, residual systematics are present, and variance from the PCA foreground cleaning is not accounted for.
@@
This provides a distribution of $\delta\hat{P}_{i}^{\rm rec}$ uncertainties on which 68th percentile limits can be taken to provide $1\sigma$ error bar estimates. Furthermore, the covariance matrix can be trivially estimated from this distribution by taking the covariance over $\hat{P}^{\rm rec}_{i}(k)\,{=}\,\hat{P}^{\rm rec}(k)\,{+}\,\delta\hat{P}^{\rm rec}_{i}(k)$. These estimates will naturally include contributions from sample variance, galaxy shot noise (from the galaxy mocks which replicate the real galaxy density), residual foregrounds and systematics, instrumental noise, and the variance associated with signal loss from eigenmode projection. Importantly, because the mocks are injected into the real data, any residual contaminants that persist in the cleaned maps are automatically included in the covariance budget.
@@
This can also be computed in the spherically- or cylindrically-averaged power spectrum space. This forms the basis for estimating the signal-to-noise of the cross-power as shown by the third panel of Figure 4. This shows the corresponding signal-to-noise distribution, highlighting which regions of $(k_{\perp},k_{\parallel})$ space are robustly measured. This directly motivated the adoption of scale cuts, illustrated by the orange lines, where modes outside the well-behaved region are excluded from the final spherically-averaged power spectrum. As the fourth panel demonstrates, these cuts remove a large chunk of the available Fourier modes. Nevertheless, this conservative strategy is necessary with the current survey volume and sensitivity. In the future, as MeerKLASS expands in area and as larger samples become available with the SKAO, many more modes will be accessible, and the statistical penalty from scale cuts will be less severe, or potentially not even needed.
@@
More robust error estimation, more conservative results: An outcome of the analysis in MeerKLASS Collaboration et al. [100] was that the estimated errors on the measured cross-power spectrum were larger than anticipated, and certainly larger than analytical methods used in the earlier intensity mapping detections [153, 56]. The mock-based transfer function method naturally captures contributions from residual systematics, the variance in signal loss due to eigenmode projection, and the presence of correlated errors between bins in Fourier space. This contrasts with earlier diagonal covariance assumptions, where only independent errors were derived based on auto-correlation noise estimation. Including these additional contributions leads to a more realistic, but inevitably more conservative, characterisation of the significance of the signal. Furthermore, the work in Chen [47], which, as discussed, revealed that the transfer function is currently complicated by mode-mixing, could also be biasing the covariance estimation. Hence, further work is planned (as discussed in Section 3.2) to investigate the numerical reconstruction routines and settle on the most robust approach for covariance estimation.
@@
As a result, the detection of the GAMA cross-correlation is quoted at the  ${\sim}\,4\sigma$ level, compared to the  ${\sim}\,7\sigma$ significance claimed for the WiggleZ field. The larger field size of WiggleZ also makes this a less direct comparison. However, it is important to stress that this more conservative result does not necessarily imply a weaker underlying signal. Instead, it reflects the fact that the new pipeline is incorporating variance at a more complete level, albeit possibly being overestimated due to mode-mixing [47]. Indeed, it was shown that if the contribution from systematics could be eliminated, driving down the covariance to the level expected for a Gaussian signal+noise (i.e. a $\chi^{2}$ consistent with uncorrelated bins and error bars representative of scatter), the GAMA-cross detection significance would increase to nearly $10\sigma$. This highlights both the robustness of the current conservative detection and the major potential gains that will be unlocked by continued improvements in calibration and systematic control.
@@ 5.3 Pursuit of a HI auto-correlation power spectrum detection
Previous single-dish intensity mapping experiments have been dominated by thermal noise at all scales, but with MeerKLASS we are now crossing the threshold where the cosmological Hi fluctuations begin to dominate the expected thermal noise on large scales with the L-band deep field observations. Figure 5 shows the predicted Hi-only auto-power spectrum from the deep field (grey dotted line), compared to an estimate of the thermal noise (red solid line) obtained from Gaussian random realisations using the known observational time within each map pixel.
On large scales ($k\,{\lesssim}\,0.15\,h{\rm Mpc}^{-1}$), the cosmological Hi power is predicted to exceed the thermal noise, and the total expected power (black dashed line) can be approximated by the sum of the two contributions. The data follow this expectation closely, with a small ( ${\sim}\,2$) mean residual factor between the model and the MeerKLASS measurements (blue data points) across all $k$. This represents the tightest upper limit to date on the Hi auto-power on large scales [a detection of a Hi auto power spectrum has been reported in 107, on smaller scales].
The MeerKLASS collaboration is pursuing several routes for improvement in future work, including splitting the data into independent sub-sets (different dishes and scan times) to isolate and mitigate systematics. A more dedicated auto-correlation study will be presented in upcoming work [101], which will build upon this first result and push towards a robust measurement of the Hi auto-power.
Figure 5: The Hi auto-power spectrum for the MeerKLASS L-band deep-field from [119]. The red solid line is the noise estimate from measuring the average power of 10 Gaussian random fields with an RMS expected to be consistent with the thermal noie level. The grey dotted line is the predicted Hi power spectrum model and the black dashed line is the combination of this Hi model plus the thermal noise estimate.
@@ 5.4 Stacking HI emission onto galaxy positions
The first demonstration of this approach with MeerKLASS was presented in MeerKLASS Collaboration et al. [100] with a more detailed follow-up and extension in Chen et al. [48]. These studies showed that by stacking the L-band deep-field maps at the positions of GAMA galaxies, a clear ($7.45\sigma$) excess signal in the stacked profile was revealed, as shown in Figure 6. However, the work also highlighted challenges that are specific to broad-beam single-dish intensity mapping: the  ${\sim}\,1^{\circ}$ beam implies that multiple galaxies fall within each resolution element, leading to double-counting and a plateau of excess emission in the stacked spectrum. Furthermore, the PCA cleaning applied to remove foregrounds introduces signal loss and spectral distortions that must be accounted for with forward modelling. Despite these complications, these studies provided the first evidence that stacking can recover meaningful information from single-dish intensity maps, complementary to power spectrum analyses.
@@
The follow-up analysis in Chen et al. [48] significantly extended this framework. Using the same MeerKLASS L-band deep-field data, the work presented a forward-modelling pipeline that incorporates clustering beyond Poisson statistics, beam convolution, and the effects of foreground cleaning. By comparing stacked images and spectra with mock simulations, they demonstrated that the central excess in the stacked signal is robust and consistent with Hi emission from the GAMA galaxies. The study also introduced improved covariance estimation based on random shuffling of galaxy positions, allowing systematic effects in the data to be characterised and constrained. Importantly, the stacked spectrum was shown to be sensitive to chromatic systematics such as primary beam ripple [15, 99], providing a new diagnostic for instrumental effects. This lays the foundation for using stacking not only as a means of measuring average Hi emission, but also as a powerful tool for validating calibration and foreground-removal pipelines.
@@ 6 Future prospects and the road ahead
21cm intensity mapping has the potential to become a valuable resource for cosmology. In the future, traditional galaxy surveys will begin to approach the limits of their reach, constrained by redshift depth and systematic effects that become increasingly dominant as statistical errors shrink. By contrast, 21cm intensity mapping can rapidly chart the three-dimensional structure of the Universe spectroscopically, uniquely reaching high redshifts and motivating next-generation experiments [129, 37, 19, 46]. The 21cm signal is inherently tied to the astrophysics of hydrogen, its abundance and distribution shaped by processes such as reionisation, star formation, and the subsequent evolution of galaxies that host neutral gas. Fortunately, going to sufficiently large, linear scales simplifies the impact of these complexities, as the connection between astrophysics and the Hi power spectrum enters primarily through the overall amplitude of the clustering [^4] Plus subdominant shot-noise terms., allowing cosmological information to be extracted with fewer degeneracies to the complex astrophysics [151, 25, 89, 28].
@@
In this section, we introduce some selected forecasts that preview the performance of the full MeerKLASS survey and provide a flavour of its potential. The formalism used in this section is detailed in Appendix A and the Fisher forecast code is publicly available at: https://github.com/meerklass/MeerFish. We will extend the complexity of these forecasts in follow-up work, but for now, we ensure that all major caveats and assumptions connected with the results are clearly mentioned. For these forecasts, we assume the 2028 MeerKLASS survey (see Table 1), that is a $10{,}000\,\text{deg}^{2}$ UHF-band survey. We assume the 2,5002{,}500\,hrs of observations have a conservative 50% loss due to RFI, hence all used observing times in the forecasts are 1,2501{,}250\,hrs. We confirm all default survey parameters assumed for MeerKLASS in Table 2, along with a next-generation SKAO-Mid Band 1 survey [128], which we also include in some forecasts.
@@ 6.1 Forecasting MeerKLASS potential
Survey sky area $A_{\mathrm{sur}}$ 10 000 deg2 20 000 deg2
Total observation time $t_{\mathrm{tot}}$ 2 500 hrs 10 000 hrs
@@
Useable observaton time $t_{\mathrm{obs}}$ 1 250 hrs 5 000 hrs
@@
Table 2: Specifications for the MeerKLASS, UHF-band, and SKAO-MID Band 1 surveys [119, 128]. 64 of the SKAO dishes will be the existing MeerKAT dishes that have $D_{\mathrm{dish}}=13.5\,\text{m}$, but for simplicity we make the approximation that all dishes have the same $15\,\text{m}$ diameter.
Figure 7: Power spectrum monopole signal-to-noise for MeerKLASS, demonstrating its competitive potential if systematics can be controlled. Current state-of-the-art DESI DR2 [64] is shown for comparison. The DESI redshift bins are used for the 21cm forecasts (indicated by the coloured points which pair to the coloured bars for each bin), plus one additional distinct high-$z$ bin for the next generation SKAO-Mid Band 1 intensity mapping survey (red point).
Figure 7 shows the forecast signal-to-noise ($S/N$) on the spherically-averaged (i.e. monopole $P_{0}$) power spectrum measurement at a nominal scale of $k\,{=}\,0.02\,h\,\text{Mpc}^{-1}$, suitable for large-scale cosmology where 21cm intensity mapping is most efficient. For simplicity, we assume in all forecasts that signal loss from foreground cleaning has been unbiasedly reconstructed. This is not an unreasonable assumption given the current development of the foreground transfer function [57, 47] (see Section 3.2), however, it is inevitable that this process will increase uncertainties on large scales (small-$k$), which, in lieu of a reliable covariance model for signal reconstruction, has not been included and we leave this extension to future work.
For comparison, Figure 7 also shows the $S/N$ from recent DESI data (coloured bars). We take the galaxy catalogue counts and survey areas from DESI DR2 [64], for each of their redshift bins, and compute the mean galaxy number densities, which are used in our modelling formalism to estimate $P/\sigma{P}$, assuming $b_{\text{g}}\,{=}\,2.0,1.2,2.1$ for the LRG, ELG and QSO samples respectively [65]. The intensity maps are matched to these DESI redshift bins where possible (indicated by the coloured points, which pair to each coloured DESI redshift bin). This suggests MeerKLASS can approximately equal the DR2 $S/N$ levels across the available UHF redshift range. We will demonstrate shortly how this can be leveraged for competitive constraints on cosmological parameters. The next generation SKAO $S/N$ is shown by the grey-dotted line, for the same redshift bins as DESI DR2, plus one additional high-redshift bin at 1.6<z< 31.6\,{<}\,z\,{<}\,3 (red point). We note that DESI will continue to gather data, and thus the DESI results presented will improve in terms of statistical precision and systematic control, leading to enhanced signal-to-noise. The 5-year DESI survey began in 2021 and is set to gather 20 million unique redshifts by its completion [61, 62].

### v2

@@ 5.2 Covariance estimation and scale cuts
The work in [MeerKLASS:2024ypg] investigated in detail how to efficiently extract the cosmological signal from MeerKLASS data and also estimate the errors on any measurements. The far-left panel of Figure 4 again shows the cross-correlation power spectrum between the MeerKLASS deep field and the GAMA galaxies but, instead of the 1D spherically-averaged power, this is cylindrically averaged into ($_{\perp}, kk_{\parallel}$) bins. This gives a more detailed view into where the signal exists in this analysis. The panel clearly shows the strong effect of the instrumental beam, where the power on small transverse scales (large $k_{\perp}$) is suppressed to nearly zero. At the opposite end, the strength of the signal at extremely small $k$ is also unreliable, which can be attributed both to survey geometry limitations and to residual contamination from foregrounds. This interpretation is supported by the transfer function analysis (shown in the second panel), which indicates that nearly all of the cosmological signal is lost ($\mathcal{T}\,{\sim}\,0$) at the very smallest $k$-modes owing to their strong degeneracy with foreground spatial structure.
@@
A major part of the investigation in [MeerKLASS:2024ypg] was to explore the use of the transfer function not only as a correction for signal loss but also as a means of estimating the covariance of the power spectrum. In earlier analyses [e.g. Cunnington:2022uzo], analytical error estimates were adopted, assuming that thermal noise dominated the error budget and that Gaussian statistics applied. However, with the improved sensitivity of MeerKLASS, these assumptions no longer hold: thermal noise is subdominant on many scales, residual systematics are present, and variance from the PCA foreground cleaning is not accounted for.
@@
This provides a distribution of $\delta\hat{P}_{i}^{\rm rec}$ uncertainties on which 68th percentile limits can be taken to provide $1\sigma$ error bar estimates. Furthermore, the covariance matrix can be trivially estimated from this distribution by taking the covariance over $\hat{P}^{\rm rec}_{i}(k)\,{=}\,\hat{P}^{\rm rec}(k)\,{+}\,\delta\hat{P}^{\rm rec}_{i}(k)$. These estimates will naturally include contributions from sample variance, galaxy shot noise (from the galaxy mocks which replicate the real galaxy density), residual foregrounds and systematics, instrumental noise, and the variance associated with signal loss from eigenmode projection. Importantly, because the mocks are injected into the real data, any residual contaminants that persist in the cleaned maps are automatically included in the covariance budget (which we discuss in more detail shortly).
@@
The covariance and error estimations can be computed in the spherical- or cylindrical-averaged power spectrum space. This forms the basis for estimating the signal-to-noise of the cross-power as shown by the third panel of Figure 4. This shows the corresponding signal-to-noise distribution, highlighting which regions of $(k_{\perp},k_{\parallel})$ space are robustly measured. This directly motivated the adoption of scale cuts, illustrated by the orange lines, where modes outside the well-behaved region are excluded from the final spherically-averaged power spectrum. As the fourth panel demonstrates, these cuts remove a large chunk of the available Fourier modes. Nevertheless, this conservative strategy is necessary with the current survey volume and sensitivity. In the future, as MeerKLASS expands in area and as larger samples become available with the SKAO, many more modes will be accessible, and the statistical penalty from scale cuts will be less severe.
Figure 5: (left): $k$-bin correlation matrix for the cross-power between the MeerKLASS L-band deep-field intensity maps and GAMA galaxies, following the semi-empirical approach that utilises the transfer function variance (detailed in the text). (right): Diagonal covariance comparison between the transfer-function based approach (blue) computed from the diagonals of the full covariance matrix, and an analytical estimation (orange), outlined in the text, used in previous analyses.
More robust error estimation, more conservative results: The normalised covariance matrix, i.e. the correlation matrix, is shown in the left panel of Figure 5 for the MeerKLASS deep-field cross power measurement. This uses the empirical transfer-function-based approach, and we clearly see non-negligible contributions from the off-diagonal elements, motivating this extension beyond the purely analytical diagonal approach in previous analyses.
An outcome of the analysis in MeerKLASS:2024ypg was that the estimated errors on the measured cross-power spectrum were larger than anticipated, and certainly larger than analytical methods used in the earlier intensity mapping detections [eBOSS:2021ebm, Cunnington:2022uzo]. We demonstrate this in the right panel of Figure 5 where we compare the (blue) diagonals of the transfer-function-based covariance (i.e. a mock-based calculation with empirical contributions from real systematics in the data) and (orange) an analytical estimation of the diagonal (Gaussian) covariance calculated from
$\color[rgb]{0,0,0}\sigma_{P_{\textrm{H{i}},\mathrm{g}}}(k)=\frac{1}{\sqrt{2N_{\mathrm{modes}}(k)}}\sqrt{\hat{P}_{\textrm{H{i}},\mathrm{g}}^{2}(k)+\hat{P}_{\textrm{H{i}}}(k)\hat{P}_{\mathrm{g}}(k)}\,.$} (17)
Here, $\hat{P}_{\textrm{H{i}}}$ and $\hat{P}_{\text{g}}$ are the estimators for the Hi and galaxy auto-correlation power spectra, inclusive of the thermal noise and galaxy shot-noise, respectively.
@@
The mock-based transfer function method naturally captures contributions from residual systematics, the variance in signal loss due to eigenmode projection, and the presence of correlated errors between bins in Fourier space. This contrasts with earlier diagonal covariance assumptions, where only independent errors were derived based on auto-correlation noise estimation (see Equation 17). Including these additional contributions leads to a more realistic, but inevitably more conservative, characterisation of the significance of the signal. Furthermore, the work in Chen:2025til, which, as discussed, revealed that the transfer function is currently complicated by mode-mixing, could also be biasing the covariance estimation. Hence, further work is planned (as discussed in Section 3.2) to investigate the numerical reconstruction routines and settle on the most robust approach for covariance estimation.
@@
As a result, the detection of the GAMA cross-correlation is quoted at the  ${\sim}\,4\sigma$ level, compared to the  ${\sim}\,7\sigma$ significance claimed for the WiggleZ field. The larger field size of WiggleZ also makes this a less direct comparison. However, it is important to stress that this more conservative result does not necessarily imply a weaker underlying signal. Instead, it reflects the fact that the new pipeline is incorporating variance at a more complete level, albeit possibly being overestimated due to mode-mixing [Chen:2025til]. Indeed, it was shown that if the contribution from systematics could be eliminated, driving down the covariance to the level expected for a Gaussian signal+noise (i.e. a $\chi^{2}$ consistent with uncorrelated bins and error bars representative of scatter), the GAMA-cross detection significance would increase to nearly $10\sigma$. This highlights both the robustness of the current conservative detection and the major potential gains that will be unlocked by continued improvements in calibration and systematic control.
@@ 5.3 Pursuit of a HI auto-correlation power spectrum detection
Previous single-dish intensity mapping experiments have been dominated by thermal noise at all scales, but with MeerKLASS we are now crossing the threshold where the cosmological Hi fluctuations begin to dominate the expected thermal noise on large scales with the L-band deep field observations.
Figure 6 shows the MeerKLASS auto-power spectrum from the deep field (blue data points) compared to a modelled prediction of a Hi-only auto-power (grey dotted line), and an estimate of the thermal noise (red solid line) obtained from Gaussian random realisations using the known observational time within each map pixel. On large scales ($k\,{\lesssim}\,0.15\,h{\rm Mpc}^{-1}$). The cosmological Hi power is predicted to exceed the thermal noise, and the total expected power (black dashed line) can be approximated by the sum of the two contributions. The data follow this expectation closely, with a small ( ${\sim}\,2$) mean residual factor between the model and the MeerKLASS measurements across all $k$. This represents the tightest upper limit to date on the Hi auto-power on large scales (a detection of a Hi auto power spectrum has been reported in Paul:2023yrr on smaller scales, and now also in CHIME:2025cee, again on smaller scales).
The MeerKLASS collaboration is pursuing several routes for improvement in future work, including splitting the data into independent sub-sets (different dishes and scan times) to isolate and mitigate systematics. A more dedicated auto-correlation study will be presented in upcoming work [MatildeAutoPk], which will build upon this first result and push towards a robust measurement of the Hi auto-power.
Figure 6: The Hi auto-power spectrum for the MeerKLASS L-band deep-field from [MeerKLASS:2024ypg]. The red solid line is the noise estimate from measuring the average power of 10 Gaussian random fields with an RMS expected to be consistent with the thermal noie level. The grey dotted line is the predicted Hi power spectrum model and the black dashed line is the combination of this Hi model plus the thermal noise estimate.
@@ 5.4 Stacking HI emission onto galaxy positions
The first demonstration of this approach with MeerKLASS was presented in MeerKLASS:2024ypg with a more detailed follow-up and extension in Chen:2025bcx. These studies showed that by stacking the L-band deep-field maps at the positions of GAMA galaxies, a clear ($7.45\sigma$) excess signal in the stacked profile was revealed, as shown in Figure 7. However, the work also highlighted challenges that are specific to broad-beam single-dish intensity mapping: the  ${\sim}\,1^{\circ}$ beam implies that multiple galaxies fall within each resolution element, leading to double-counting and a plateau of excess emission in the stacked spectrum. Furthermore, the PCA cleaning applied to remove foregrounds introduces signal loss and spectral distortions that must be accounted for with forward modelling. Despite these complications, these studies provided the first evidence that stacking can recover meaningful information from single-dish intensity maps, complementary to power spectrum analyses.
@@
The follow-up analysis in Chen:2025bcx significantly extended this framework. Using the same MeerKLASS L-band deep-field data, the work presented a forward-modelling pipeline that incorporates clustering beyond Poisson statistics, beam convolution, and the effects of foreground cleaning. By comparing stacked images and spectra with mock simulations, they demonstrated that the central excess in the stacked signal is robust and consistent with Hi emission from the GAMA galaxies. The study also introduced improved covariance estimation based on random shuffling of galaxy positions, allowing systematic effects in the data to be characterised and constrained. Importantly, the stacked spectrum was shown to be sensitive to chromatic systematics such as primary beam ripple [2021MNRAS.502.2970A, Matshawule:2020fjz], providing a new diagnostic for instrumental effects. This lays the foundation for using stacking not only as a means of measuring average Hi emission, but also as a powerful tool for validating calibration and foreground-removal pipelines.
@@ 6 Future prospects and the road ahead
21cm intensity mapping has the potential to become a valuable resource for cosmology. In the future, traditional galaxy surveys will begin to approach the limits of their reach, constrained by redshift depth and systematic effects that become increasingly dominant as statistical errors shrink. By contrast, 21cm intensity mapping can rapidly chart the three-dimensional structure of the Universe spectroscopically, uniquely reaching high redshifts and motivating next-generation experiments [PUMA:2019jwd, Burns:2021ndk, 2023arXiv230110345B, Chen:2024tvn]. The 21cm signal is inherently tied to the astrophysics of hydrogen, its abundance and distribution shaped by processes such as reionisation, star formation, and the subsequent evolution of galaxies that host neutral gas. Fortunately, going to sufficiently large, linear scales simplifies the impact of these complexities, as the connection between astrophysics and the Hi power spectrum enters primarily through the overall amplitude of the clustering [^5] Plus subdominant shot-noise terms., allowing cosmological information to be extracted with fewer degeneracies with the complex astrophysics [Wolz:2017rlw, Beane:2018pmx, Kovetz:2019uss, Bernal:2019jdo].
@@
In this section, we introduce some selected forecasts that preview the performance of the full MeerKLASS survey and provide a flavour of its potential. The formalism used in this section is detailed in Appendix A and the Fisher forecast code is publicly available at: https://github.com/meerklass/MeerFish. We will extend the complexity of these forecasts in follow-up work, but for now, we ensure that all major caveats and assumptions connected with the results are clearly mentioned. For these forecasts, we assume the 2028 MeerKLASS survey (see Table 1), that is a $10{,}000\,\text{deg}^{2}$ UHF-band survey. We assume the 2,5002{,}500\,hrs of observations have a conservative 50% loss due to RFI, hence all used observing times in the forecasts are 1,2501{,}250\,hrs. We confirm all default survey parameters assumed for MeerKLASS in Table 2, along with a next-generation SKAO-Mid Band 1 survey [SKA:2018ckk], which we also include in some forecasts.
@@ 6.1 Forecasting MeerKLASS potential
Survey sky area $A_{\mathrm{sur}}$ 10,000 deg2 20,000 deg2
Total observation time $t_{\mathrm{tot}}$ 2,500 hrs 10,000 hrs
@@
Useable observaton time $t_{\mathrm{obs}}$ 1,250 hrs 5,000 hrs
@@
Table 2: Specifications for the MeerKLASS, UHF-band, and SKAO-MID Band 1 surveys [MeerKLASS:2017vgf, SKA:2018ckk]. 64 of the SKAO dishes will be the existing MeerKAT dishes that have $D_{\mathrm{dish}}=13.5\,\text{m}$, but for simplicity we make the approximation that all dishes have the same $15\,\text{m}$ diameter.
Figure 8: Power spectrum monopole signal-to-noise for MeerKLASS, demonstrating its competitive potential if systematics can be controlled. Current state-of-the-art DESI DR2 [DESI:2025zgx] is shown for comparison. The DESI redshift bins are used for the 21cm forecasts (indicated by the coloured points which pair to the coloured bars for each bin), plus one additional distinct high-$z$ bin for the next generation SKAO-Mid Band 1 intensity mapping survey (red point).
Figure 8 shows the forecast signal-to-noise ($S/N$) on the spherically-averaged (i.e. monopole $P_{0}$) power spectrum measurement at a nominal scale of $k\,{=}\,0.02\,h\,\text{Mpc}^{-1}$, suitable for large-scale cosmology where 21cm intensity mapping is most efficient. For simplicity, we assume in all forecasts that signal loss from foreground cleaning has been unbiasedly reconstructed. This is not an unreasonable assumption given the current development of the foreground transfer function [Cunnington:2023jpq, Chen:2025til] (see Section 3.2), however, it is inevitable that this process will increase uncertainties on large scales (small-$k$), which, in lieu of a reliable covariance model for signal reconstruction, has not been included and we leave this extension to future work.
For comparison, Figure 8 also shows the $S/N$ from recent DESI data (coloured bars). We take the galaxy catalogue counts and survey areas from DESI DR2 [DESI:2025zgx], for each of their redshift bins, and compute the mean galaxy number densities, which are used in our modelling formalism to estimate $P/\sigma{P}$, assuming $b_{\text{g}}\,{=}\,2.0,1.2,2.1$ for the LRG, ELG and QSO samples respectively [DESI:2025qqy]. The intensity maps are matched to these DESI redshift bins where possible (indicated by the coloured points, which pair to each coloured DESI redshift bin). This suggests MeerKLASS can approximately equal the DR2 $S/N$ levels across the available UHF redshift range. We will demonstrate shortly how this can be leveraged for competitive constraints on cosmological parameters. The next generation SKAO $S/N$ is shown by the grey-dotted line, for the same redshift bins as DESI DR2, plus one additional high-redshift bin at 1.6<z< 31.6\,{<}\,z\,{<}\,3 (red point). We note that DESI will continue to gather data, and thus the DESI results presented will improve in terms of statistical precision and systematic control, leading to enhanced signal-to-noise. The 5-year DESI survey began in 2021 and is set to gather 20 million unique redshifts by its completion [DESI:2025fxa, DESI:2024uvr].